OpenStudy (anonymous):

A small toy company makes two types of toys: dolls and toy cars. The expense to make a toy car is $5. The expense to make a doll is $7. The company spends $780 a day making toys. They also make 20 toy cars fewer than twice the number of dolls they make. Construct a system of equations that describes this situation.

OpenStudy (anonymous):

@Hero needs to help.

OpenStudy (anonymous):

Or @jackellyn (:

OpenStudy (jackellyn):

let c be toy cars d be dolls The expense to make a toy car is $5. The expense to make a doll is $7. The company spends $780 a day making toys 780 = -5c -7d They also make 20 toy cars fewer than twice the number of dolls they make: c=2d-20 So the system of equations is: 780 = -5c -7d c=2d-20

OpenStudy (anonymous):

Is that answer right @Loujoelou

OpenStudy (anonymous):

I'm just wondering, @jackellyn why wouldn't the first equation be 5c+7t=780?

OpenStudy (jackellyn):

@Loujoelou , minor error I meant to put -780 = -5c - 7d which is equivalent to 780 = 5c +7d because they are losing money in all three aspects (spend, expense, expense).